OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line?

OpenStudy (ankit042):

m =1 as slope is 45 y-y1 = m(x-x1) (x1,y1) = (0,1)

OpenStudy (anonymous):

thx

OpenStudy (ankit042):

np :)